# Ngoc BuiY , Shubham SharmaJ, Simran LambaJ, Saumitra MishraJ, Rex YingY

### v1

@@ 4.2 Training
Here, $\mathcal{D}_{\mathrm{KL}}$ is the standard forward Kullback-Leibler divergence (kullback1951information).
@@
$\mathcal{L}_{\mathrm{cap}}=\frac{1}{T(T-M)}\sum_{t=1}^{T}\max~\{0,\sum_{i=1}^{t}\beta_{i}^{t-i}-M\},$ (5)
@@
Hardware-aware Computation. Retention-gated attention is fully parallelizable and compatible with FlashAttention-style kernels (dao2023flashattention). We implement it with FlexAttention (dong2024flex) plus a custom Triton kernel for the capacity loss $\mathcal{L}_{\mathrm{cap}}$, performing forward/backward without materializing the full attention or $\beta$ matrices. This enables long-context training (up to 128K tokens on four H100 GPUs) with minor overhead versus standard parameter-efficient fine-tuning.
@@ 4.3 Inference
Complexity. Our inference is simpler and more efficient than existing works, including pure heuristic baselines (li2024snapkv). Throughput and runtime comparisons are in Appendix A.2.
@@ 5.1 Long Generation Evaluation
Benchmarks. Following prior work (gao2025seerattention; cai2025r), we evaluate on standard math-reasoning suites—AIME24 (aime2024), GSM8K (cobbe2021training), and MATH-500 (hendrycks2021measuring). To assess performance beyond math reasoning and under long-context settings, we also report results on LongProc (ye2025longproc). Following (gao2025seerattention), we report average pass@1 accuracy over 64 samples for AIME24 and 8 samples for GSM8K, MATH-500. We use greedy decoding for LongProc as the default in the benchmark.
Base Models. Following (gao2025seerattention), we mainly use Qwen3’s family models (yang2025qwen3), including Qwen3-1.7B, Qwen3-4B, Qwen3-8B, Qwen3-14B and DeepSeek R1 Distill (guo2025deepseek) variants including, DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Lllam-8B. We report the results with Qwen3 models in the main paper, and the remaining is in Appendix B.
Baselines. We compare our method against SeerAttn-R (gao2025seerattention), R-KV (cai2025r), SnapKV (li2024snapkv), H2O (zhang2023h2o), StreamingLLM (xiao2023efficient). R-KV, SnapKV, H2O, and StreamingLLM are heuristic, recency-driven KV eviction policies for long-form generation under a fixed memory budget. SeerAttn-R is a learnable KV retrieval approach for reasoning tasks: rather than evicting, it offloads the full KV cache to host memory and uses recent queries to fetch relevant blocks for attention. KV retrieval methods preserve all past information but require nontrivial CPU–GPU orchestration and incur offloading overhead. We therefore treat SeerAttn-R as a strong learnable baseline, and R-KV/SnapKV as representative eviction baselines.
Implementation Details. We train the retention gates using OpenR1-MATH-220k (openr1math2025) dataset, similar to (gao2025seerattention). Note that we only train the retention gates’ weights while keeping the original model parameters frozen. We set the objective hyperparameters $\lambda_{\mathrm{cap}}=1.0$ and the memory capacity $M=512$. Each transformer block has a retention gate $g$, which is a single MLP layer with the hidden dimension of $512$, thus having dimensions of $d\to 512\to h$, where $h$ is the number of KV heads. We use the activation function as the default activation function in MLP layers of the base model. We initialize the bias in the retention gates to a large value (e.g., $b=8.0$) to begin training with minimal forgetting or compression. All trainings are on 4 H100 80G GPUs.
@@ 5.1.1 Quantitative Result
Math Reasoning Tasks. Figure 3 shows our method outperforming all baselines by a large margin, especially in low-budget regimes. Notably, TRIM-KV surpasses attention-guided methods (R-KV, SnapKV) even when they are given $4\times$ KV budget. Under the same budget, i.e.1024 for AIME24 and 128 for GSM8K/MATH-500, it yields a 198% relative improvement over these baselines. Against the SOTA learnable KV retrieval baseline, TRIM-KV outperforms SeerAttn-R across all settings, yielding a 58.4% pass@1 gain at the same budget. Crucially, TRIM-KV operates in a pure KV-eviction regime, a stricter setting than the KV retrieval methods such as SeerAttn-R, and thus avoids CPU–GPU offloading overhead. In some settings, like for Qwen3-4B model and AIME24 dataset, our method can even surpass the standard full KV cache. These results suggest that a large fraction of KV-cache tokens in reasoning models is redundant and can be discarded without degrading performance.
@@ 5.1.2 Qualitative Result
To examine the eviction policy learned by our retention gates, we run TRIM-KV on Qwen3-4B for the first example in AIME24 (see Figure 14 for a visualization of the example). Figure 5ab show the mean retention score, averaged over all layers and heads, for each token in the example sequence. Aligning with our intuition, retention gates assign high scores to task-relevant tokens (e.g., ometer, shop, walk, minutes) and to the initial token <|im_start|>, which often serves as an attention sink. In contrast, whitespace and punctuation receive low retention scores and are discarded early, yielding short lifespans in the KV cache. Next, we examine retention scores and eviction decisions at layer–head granularity.
@@
Emergent Eviction Heuristics. Figure 4 visualizes the retention scores $\beta^{t-i}_{i}$ and eviction decisions $\alpha_{ti}$ for selected layers and heads. Many eviction heuristics, such as attention sinks (xiao2023efficient), sliding windows (zhu2021long), A-shape (jiang2024minference), emerge naturally from our learned policy without being hard-coded, and they adapt to the functional roles of individual layers and heads. For instance, sliding-window behavior is more common in early layers, whereas attention sinks appear more frequently in later layers (see Figure 12 and 13 for a comprehensive view). Moreover, TRIM-KV adapts the window size by layer and head: in Layer 1/Head 1, tokens receive nearly uniform retention scores, so the KV cache behaves like a recency-biased window that keeps the most recent tokens; in Layer 0/Head 0, multiple sliding windows of varying widths emerge from the learned policy; in Layer 15/Head 2, no sliding window is observed because certain tokens receive substantially higher retention than others, suggesting a specialized functional role for this head. The A-shaped pattern typically appears in layers that emphasize instruction/problem-statement tokens (e.g., Layer 20/Head 5 and Layer 30/Head 4) or chain-of-thought/reasoning prompts (e.g., Layer 35/Head 1). These heads also exhibit context switching, where small, dense lower-triangular blocks emerge and then fade quickly when the context changes or a sentence completes. To the best of our knowledge, the absence of a sliding window, the presence of multiple coexisting windows, and context switching are newly observed eviction patterns that arise naturally from our learned policy.
@@
Token Retention Enables Interpretability. Beyond guiding eviction policy, token-level retention scores provide a diagnostic tool for analyzing the functional roles of individual KV heads in the base LLM. Visualizing retention scores alongside the tokens preserved in the KV cache after generation reveals distinct specializations: some heads emphasize a recency window (Figure 17), whereas others preferentially retain mathematical tokens-numbers and operators (Figures 14 and 20)-or variables (Figure 23), as well as problem-description tokens (Figures 19 and 25) and chain-of-thought instructions (Figure 26). Even function or filler words, such as pronouns, prepositions, conjunctions, wait and let, tend to be kept by specific heads (Figures 18 and 24). In particular, heads that exhibit context-switching patterns (e.g., Layer 30/Head 4 and Layer 16/Head 6) tend to retain the period token while discarding others (Figures 22 and 21). We hypothesize that, in these heads, periods act as implicit gist tokens (mu2023learning), summarizing the information in the preceding sentences.
@@
Our analyses indicate that KV heads in LLM develop different functional roles and therefore keep different types of tokens. These tokens are often dispersed across the context rather than forming contiguous chunks, as each already captures contextual information. This observation contrasts with existing approaches (yuan2025native; gao2025seerattention) that advocate chunk- or block-based KV-cache. Instead, we show that keeping a small number of high-context tokens is more budget-effective.
Budget Allocation. Figure 5c reports head- and layer-wise sparsity estimated from the retention scores i.e., $1-\frac{2}{T(T+1)}\sum_{i<t}\beta_{i}^{\,t-i}$. We observe that later layers are typically sparser than earlier ones, consistent with prior findings in (cai2024pyramidkv). Practically, the retention scores enable heterogeneous budgets across KV heads under a global constraint by evicting tokens with low global retention. However, existing KV-cache and FlashAttention implementations assume uniform sequence lengths across heads within a layer; enabling efficient per-head variable-length caches is left to future work.
5.2 Long-Context Evaluation
To further stress-test our approach on long-context tasks, we evaluate it on LongMemEvalS (wu2024longmemeval), a benchmark for assessing chat assistants’ long-term interactive memory (contexts up to 123K tokens with Qwen3) and SCBench (li2024scbench), a benchmark adapted from RULER (hsieh2024ruler) and InfiniteBench (zhang2024bench) to evaluate long-context understanding of KV cache compression methods under the same context but different queries. We use Qwen3-4B-Instruct (qwen3_4b_instruct) as the base model and train the retention gates on SynthLong (cerebras2025synthlong), BookSum (kryscinski2021booksum), and Buddhi (singhal2024buddhi), whose context lengths extend up to 128K tokens. All other experimental settings follow Section 5.1.
@@ 5.2 Long-Context Evaluation
As shown in Table 3 and Table 2 show the results on LongMemEval and SCBench benchmarks respectively. On LongMemEval, TRIM-KV outperforms baselines by a significant margin. Especially, TRIM-KV can maintain the performance of a full cache while using just $25\%$ of the KV budget, while other baselines drop sharply. On SCBench, TRIM-KV also shows a competitive performance across most of the evaluated tasks. On challenging retrieval tasks such as Retr.KV and Code.RepoQA, all KV eviction methods failed to give a reasonable performance because the context is incompressible. This is already observed in (li2024scbench). These results highlight our method’s advantage on both long-context and long-generation tasks, whereas most prior work targets either the prefill or the generation stage, but not both. More details and results for this experiment and LongBench (bai2025longbench) are provided in Appendix B.
@@ 5.3 Ablation Studies
(TRIM-KV $\mathcal{L}_{\mathrm{KL}}$)4096 71.4
(TRIM-KV $\mathcal{L}_{\mathrm{NTP}}$)4096 70.7
@@ Appendix A Methodology
2 $\mathbf{q}_{t}\leftarrow\mathbf{W}_{Q}\mathbf{x}_{t}$; $\mathbf{k}_{t}\leftarrow\mathbf{W}_{K}\mathbf{x}_{t}$; $\mathbf{v}_{t}\leftarrow\mathbf{W}_{V}\mathbf{x}_{t}$; $\beta_{t}=g(\mathbf{x}_{t})$;
@@
3 $\mathbf{K}_{t}\leftarrow\mathbf{K}_{t-1}\cup\mathbf{k}_{t}$; $\mathbf{V}_{t}\leftarrow\mathbf{V}_{t-1}\cup\mathbf{v}_{t}$; $\mathbf{B}_{t}\leftarrow\mathbf{B}_{t-1}\cup\beta_{t}$; $S_{t}\leftarrow S_{t-1}\cup\{t\}$;
@@
4 $\mathbf{o}_{t}\leftarrow\textsc{FlashAttn}(\mathbf{q}_{t},\mathbf{K}_{t},\mathbf{V}_{t})$;
@@
5 while $|S_{t}|>M$ do
6  $j_{\mathrm{evic}}\leftarrow\arg\min\ \{\beta_{j}^{t-j}|j\in S_{t}\}$;
7  Remove $\mathbf{K}_{t}[j_{\mathrm{evic}}]$, $\mathbf{V}_{t}[j_{\mathrm{evic}}]$, $\mathbf{B}_{t}[{j_{\mathrm{evic}}}]$;
8  $S_{t}\leftarrow S_{t}\setminus\{j_{\mathrm{evic}}\}$;
9
10 end while
@@ A.2 Complexity
Memory efficiency. Like other KV-eviction schemes, TRIM-KV uses a fixed-size cache with $\mathcal{O}(M)$ slots, independent of sequence length $T$. For each token (per head), it stores a single scalar retention score $\beta_{i}$, adding only $\approx 1/d_{h}$ overhead, where $d_{h}$ is the dimension of the key and vector states, relative to the KV states, which is negligible in practice. Unlike R-KV [cai2025r], TRIM-KV does not store queries.

### v2

@@ 4.2 Training
Here, $\mathcal{D}_{\mathrm{KL}}$ is the standard forward Kullback-Leibler divergence (Kullback and Leibler, 1951).
@@
$\mathcal{L}_{\mathrm{cap}}=\frac{1}{T}\sum_{t=1}^{T}\frac{1}{t}\max~\{0,\sum_{i=1}^{t}\beta_{i}^{t-i}-M\},$ (5)
@@
Hardware-aware Computation. Retention-gated attention is fully parallelizable and compatible with FlashAttention-style kernels (Dao, 2023). We implement it with FlexAttention (Dong et al., 2024) plus a custom Triton kernel for the capacity loss $\mathcal{L}_{\mathrm{cap}}$, performing forward/backward without materializing the full attention or $\beta$ matrices. This enables long-context training (up to 128K tokens on four H100 GPUs) with minor overhead versus standard parameter-efficient fine-tuning.
@@ 4.3 Inference
Complexity. Our inference is simpler and more efficient than existing works, including pure heuristic baselines (Li et al., 2024c). Table 6 shows that, at 32K context, TRIM-KV achieves $\sim\!2\times$ higher decoding throughput than full-cache decoding and even faster than SnapKV, a purely heuristic method. More details are provided in Appendix A.2.
@@ 5.1 Long Generation Evaluation
Benchmarks. Following prior work (Gao et al., 2025; Cai et al., 2025), we evaluate on standard math-reasoning suites—AIME24 (Art of Problem Solving, 2024), GSM8K (Cobbe et al., 2021), and MATH-500 (Hendrycks et al., 2021). To assess performance beyond math reasoning and under long-context settings, we also report results on LongProc (Ye et al., 2025). Following (Gao et al., 2025), we report average pass@1 accuracy over 64 samples for AIME24 and 8 samples for GSM8K, MATH-500. We use greedy decoding for LongProc as the default in the benchmark.
Base Models. Following (Gao et al., 2025), we mainly use Qwen3’s family models (Yang et al., 2025), including Qwen3-1.7B, Qwen3-4B, Qwen3-8B, Qwen3-14B and DeepSeek R1 Distill (Guo et al., 2025) variants including, DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Lllam-8B. We report the results with Qwen3 models in the main paper, and the remaining is in Appendix B.
Baselines. We compare our method against SeerAttn-R (Gao et al., 2025), R-KV (Cai et al., 2025), SnapKV (Li et al., 2024c), H2O (Zhang et al., 2023), StreamingLLM (Xiao et al., 2023). R-KV, SnapKV, H2O, and StreamingLLM are heuristic, recency-driven KV eviction policies for long-form generation under a fixed memory budget. SeerAttn-R is a learnable KV retrieval approach for reasoning tasks: rather than evicting, it offloads the full KV cache to host memory and uses recent queries to fetch relevant blocks for attention. KV retrieval methods preserve all past information but require nontrivial CPU–GPU orchestration and incur offloading overhead. We therefore treat SeerAttn-R as a strong learnable baseline, and R-KV/SnapKV as representative eviction baselines.
Implementation Details. We train the retention gates using OpenR1-MATH-220k (Hugging Face, 2025) dataset, similar to (Gao et al., 2025). Note that we only train the retention gates’ weights while keeping the original model parameters frozen. We set the objective hyperparameters $\lambda_{\mathrm{cap}}=1.0$ and the memory capacity $M=256$. Each transformer block has a retention gate $g$, which is a single MLP layer with the hidden dimension of $512$, thus having dimensions of $d\to 512\to h$, where $h$ is the number of KV heads. We use the activation function as the default activation function in MLP layers of the base model. We initialize the bias in the retention gates to a large value (e.g., $b=18.0$) to begin training with minimal forgetting or compression. All trainings are on 4 H100 80G GPUs.
@@ 5.1.1 Quantitative Result
Math Reasoning Tasks. Figure 3 shows our method outperforming all baselines by a large margin, especially in low-budget regimes. Notably, TRIM-KV surpasses attention-guided methods (R-KV, SnapKV) even when they are given $4\times$ KV budget. Under the same budget, i.e.512 for AIME24 and 128 for GSM8K/MATH-500, it yields a 198.4% relative improvement over these baselines. Against the SOTA learnable KV retrieval baseline, TRIM-KV outperforms SeerAttn-R across all settings, yielding a 58.9% pass@1 gain at the same budget. Crucially, TRIM-KV operates in a pure KV-eviction regime, a stricter setting than the KV retrieval methods such as SeerAttn-R, and thus avoids CPU–GPU offloading overhead. In some settings, like for Qwen3-4B model and AIME24 dataset, our method can even surpass the standard full KV cache. These results suggest that a large fraction of KV-cache tokens in reasoning models is redundant and can be discarded without degrading performance.
@@ 5.1.2 Qualitative Result
To examine the eviction policy learned by our retention gates, we run TRIM-KV on Qwen3-4B for the first example in AIME24. Figure 5ab show the mean retention score, averaged over all layers and heads, for each token in the example sequence. Aligning with our intuition, retention gates assign high scores to task-relevant tokens (e.g., ometer, shop, walk, minutes) and to the initial token <|im_start|>, which often serves as an attention sink. In contrast, whitespace and punctuation receive low retention scores and are discarded early, yielding short lifespans in the KV cache. Next, we examine retention scores and eviction decisions at layer–head granularity.
@@
Emergent Eviction Heuristics. Figure 4 visualizes the retention scores $\beta^{t-i}_{i}$ and eviction decisions $\alpha_{ti}$ for selected layers and heads. Many eviction heuristics, such as attention sinks (Xiao et al., 2023), sliding windows (Zhu et al., 2021), A-shape (Jiang et al., 2024), emerge naturally from our learned policy without being hard-coded, and they adapt to the functional roles of individual layers and heads. For instance, sliding-window behavior is more common in early layers, whereas attention sinks appear more frequently in later layers (see Figure 11 and 12 for a comprehensive view). Moreover, TRIM-KV adapts the window size by layer and head: in Layer 1/Head 1, tokens receive nearly uniform retention scores, so the KV cache behaves like a recency-biased window that keeps the most recent tokens; in Layer 0/Head 0, multiple sliding windows of varying widths emerge from the learned policy; in Layer 15/Head 2, no sliding window is observed because certain tokens receive substantially higher retention than others, suggesting a specialized functional role for this head. The A-shaped pattern typically appears in layers that emphasize instruction/problem-statement tokens (e.g., Layer 20/Head 5 and Layer 30/Head 4) or chain-of-thought/reasoning prompts (e.g., Layer 35/Head 1). These heads also exhibit context switching, where small, dense lower-triangular blocks emerge and then fade quickly when the context changes or a sentence completes. To the best of our knowledge, the absence of a sliding window, the presence of multiple coexisting windows, and context switching are newly observed eviction patterns that arise naturally from our learned policy.
@@
Token Retention Enables Interpretability. Beyond guiding eviction policy, token-level retention scores provide a diagnostic tool for analyzing the functional roles of individual KV heads in the base LLM. Visualizing retention scores alongside the tokens preserved in the KV cache after generation reveals distinct specializations: some heads emphasize a recency window (Figure 14 and 17), whereas others preferentially retain mathematical tokens-numbers and operators (Figures 19), as well as problem-description tokens (Figures 16). Even function or filler words, such as pronouns, prepositions, conjunctions, wait and let, tend to be kept by specific heads (Figures 13). In particular, some heads tend to retain the period token while discarding others (Figures 18). We hypothesize that, in these heads, periods act as implicit gist tokens (Mu et al., 2023), summarizing the information in the preceding sentences.
@@
Our analyses indicate that KV heads in LLM develop different functional roles and therefore keep different types of tokens. These tokens are often dispersed across the context rather than forming contiguous chunks, as each already captures contextual information. This observation contrasts with existing approaches (Yuan et al., 2025; Gao et al., 2025) that advocate chunk- or block-based KV-cache. Instead, we show that keeping a small number of high-context tokens is more budget-effective.
Budget Allocation. Figure 5c reports head- and layer-wise sparsity estimated from the retention scores i.e., $1-\frac{2}{T(T+1)}\sum_{i<t}\beta_{i}^{\,t-i}$. We observe that later layers are typically sparser than earlier ones, consistent with prior findings in (Cai et al., 2024). Practically, the retention scores enable heterogeneous budgets across KV heads under a global constraint by evicting tokens with low global retention. However, existing KV-cache and FlashAttention implementations assume uniform sequence lengths across heads within a layer; enabling efficient per-head variable-length caches is left to future work.
5.2 Long-Context Evaluation
Long-context Decoding. We evaluate TRIM-KV on LongMemEvalS (Wu et al., 2024a) and SCBench (Li et al., 2024b), two 128K-token benchmarks for long-term memory and long-context KV compression. We use Qwen3-4B-Instruct (Qwen3, 2025) as the base model and train the retention gates on SynthLong (Lazarevich et al., 2025), BookSum (Kryściński et al., 2021), and Buddhi (Singhal, 2024) with $M=1024$. All other experimental settings follow Section 5.1. As shown in Table 3 and Table 2, TRIM-KV significantly outperforms baselines on LongMemEval, maintaining a good performance with only 25% of the budget, while others degrade sharply. On SCBench, it remains competitive across most tasks, though all KV eviction methods struggle on incompressible retrieval tasks (e.g., Retr.KV, Code.RepoQA), consistent with (Li et al., 2024b). Overall, TRIM-KV shows strong performance in both long-context and long-generation settings, unlike prior work that targets only one stage. More details of this experiment are in Appendix B.2.
@@ 5.2 Long-Context Evaluation
Chunk Prefilling. We follow Huang et al. (2024) to evaluate TRIM-KV on the chunk-prefill settings. We train Phi3-mini-128K (Abdin et al., 2024) on LongAlpaca (Chen et al., 2023) with $M=512$. All experimental settings follow (Huang et al., 2024) for fair comparison, detailed in Section B.3. Table 10 shows that TRIM-KV outperforms the LocRet, a trainable KV eviction baseline, and even full KV inference, by a significant margin.
Method Acc $\Delta$ (%)
[…]
LocRet 28.03 -2.64
[…]
Table 4: Results on LongBench-V2.
@@ 5.3 Ablation Studies
(TRIM-KV $\mathcal{L}_{\mathrm{KL}}$)4096 72.1
(TRIM-KV $\mathcal{L}_{\mathrm{NTP}}$)4096 72.5
@@ 6 Conclusion and Future Work
Disclaimer
This paper was prepared for informational purposes by the Artificial Intelligence Research group of JPMorgan Chase & Co. and its affiliates ("JP Morgan") and is not a product of the Research Department of JP Morgan. JP Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
@@ Appendix A Methodology
$\mathbf{q}_{t}\leftarrow\mathbf{W}_{Q}\mathbf{x}_{t}$; $\mathbf{k}_{t}\leftarrow\mathbf{W}_{K}\mathbf{x}_{t}$; $\mathbf{v}_{t}\leftarrow\mathbf{W}_{V}\mathbf{x}_{t}$; $\beta_{t}=g(\mathbf{x}_{t})$;
@@
$\mathbf{K}_{t}\leftarrow\mathbf{K}_{t-1}\cup\mathbf{k}_{t}$; $\mathbf{V}_{t}\leftarrow\mathbf{V}_{t-1}\cup\mathbf{v}_{t}$; $\mathbf{B}_{t}\leftarrow\mathbf{B}_{t-1}\cup\beta_{t}$; $S_{t}\leftarrow S_{t-1}\cup\{t\}$;
@@
$\mathbf{o}_{t}\leftarrow\textsc{FlashAttn}(\mathbf{q}_{t},\mathbf{K}_{t},\mathbf{V}_{t})$;
@@
while $|S_{t}|>M$ do
  $j_{\mathrm{evic}}\leftarrow\arg\min\ \{\beta_{j}^{t-j}|j\in S_{t}\}$;
  Remove $\mathbf{K}_{t}[j_{\mathrm{evic}}]$, $\mathbf{V}_{t}[j_{\mathrm{evic}}]$, $\mathbf{B}_{t}[{j_{\mathrm{evic}}}]$;
  $S_{t}\leftarrow S_{t}\setminus\{j_{\mathrm{evic}}\}$;
end while
@@ A.2 Complexity
Memory efficiency. Like other KV-eviction schemes, TRIM-KV uses a fixed-size cache with $\mathcal{O}(M)$ slots, independent of sequence length $T$. For each token (per head), it stores a single scalar retention score $\beta_{i}$, adding only $\approx 1/d_{h}$ overhead, where $d_{h}$ is the dimension of the key and vector states, relative to the KV states, which is negligible in practice. Unlike R-KV [Cai et al., 2025], TRIM-KV does not store queries.